# Towards a Faithful Quantumness Certification Functional for One-Dimensional Continuous-Variable Systems

### v1

@@ Abstract
If the phase space-based Sudarshan-Glauber distribution, $P_{{{\varrho}}}$, has negative values the quantum state, $\varrho$}}, it describes is nonclassical. Due to $P$’s singular behaviour this simple criterion is impractical to use. Recent work [Bohmann and Agudelo, Phys. Rev. Lett. 124, 133601 (2020)] presented a general, sensitive, and noise-tolerant certification functional, $\xi_{P}$, for the detection of non-classical behaviour of quantum states $P_{{{\varrho}}}$. There, it was shown that when this functional takes on negative values somewhere in phase space, $\xi_{P}(x,p)<0$, this is sufficient to certify the nonclassicality of a state. Here we give examples where this certification fails. We investigate states which are known to be nonclassical but the certification functions is positive $\xi(x,p)\geq 0$ everywhere in phase space. We generalize $\xi$ giving it an appealing form which allows for improved certification. This way we generate the best family of certification functions available so far. Yet, they also fail for very weakly nonclassical states, in other words, the question how to faithfully certify quantumness remains an open question.
I Introduction
It is desirable to be able to discriminate between quantum states on the one hand and classical states on the other. So far no simple approach that works for theoretical studies and experimentally reconstructed states alike has been found to perform this discrimination faithfully. All known approaches to discrimination are either unworkable, or they do not work unambiguously, or both.
Here we generalize a recently devised quantumness certification functional, $\xi$ [1], which constitutes the most sensitive approach so far. Our generalization provides insight into the structure of $\xi$ and makes it more sensitive still, yet, the goal of a universally faithful approach to quantum versus classical discrimination remains elusive.
In the 1960s it was established that the Sudarshan-Glauber phase space distribution, $P$ [2, 3, 4, 5, 6, 7, 8], should be used for the very definition of nonclassical behaviour [4, 5, 1]. If a state, $\varrho$}}, has a distribution $P_{{{\varrho}}}(x,p)$ with negative values somewhere at position $x$ and momentum $p$ in (cartesian) phase space  [^1] Instead of using complex coherent amplitudes to parametrize phase space we follow the convention for (quantum-mechanical) phase space of Ref. [8] using position $x$ and momentum $p$., then that state is nonclassical.
@@ I Introduction
Very useful alternatives to Sudarshan-Glauber’s $P$ distribution are Wigner’s, $W$ [13], and Husimi’s, $Q$ [14], distributions. Both are more well-behaved functions than $P$ is, but at the expense of not being able to certify that a state is nonclassical.
This situation has led to many different attempts to devise a reliable strategy for certification of the nonclassical nature of a state (see [15, 16, 17, 18, 19, 1, 20, 21, 22] and the many references therein). Similarly, a plethora of measures of nonclassicality [23, 24, 25, 26, 27, 15, 28, 29, 30, 31, 32, 33, 34, 16, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 21, 47] have been devised. None of these approaches faithfully discriminate between classical and quantum states: faithful discrimination should always detect and certify that a state is nonclassical, and it should certify only classical states as classical (for phase space-based approaches it should show that negative values are absent). We will from now on refer to this elusive ability as being ‘faithfully discriminating’.
Quantumness certification approaches devised so far, typically, mislabel states that have quantum features as classical. In the context of phase space-based approaches, this is often due to the fact that mapping functions from the potentially very singular $P$-distributions to better behaved ones, such as Wigner and Husimi distributions, washes out the very features we intend to detect.
Here we consider the recently devised phase space-based certification functional for single mode systems $\xi_{{{\varrho}}}(x,p)$, introduced by Bohmann and Agudelo in Ref. [1], in which they showed that all classical states obey $\xi_{{{\varrho}}}(x,p)\geq 0$. This is to be understood as the statement that the occurrence of negative values $\xi_{{{\varrho}}}(x,p)<0$ somewhere in phase space $(x,p)$ is sufficient to show that a state is nonclassical [1]. What is missing for $\xi$ of [1], however, is necessity: to be faithfully discriminating it would be necessary that whenever a state is nonclassical we find that $\xi(x,p)<0$ in phase space somewhere. Below, we will show that there are states $\rho$ which are (weakly) nonclassical and yet for these states $\xi_{\rho}\geq 0$, across phase space.
$\xi_{{{\varrho}}}(x,p)$ is a functional of fixed form independent of context such as the type of system or the type of state the system is in. $\xi_{{{\varrho}}}$ applies to all states of one-dimensional continuous-variable systems, irrespective of whether the state is pure or mixed and whether it has small, large, or even macroscopic [48] extent, irrespective of context, hamiltonian, environment or measurement procedures [49], in this sense $\xi$ is universal, meaning it is universally applicable.
@@
II Smeared phase space distributions
Smeared phase space distributions, ${\tilde{\cal P}}(S)$, expressed in terms of a reference distribution ${\tilde{\cal P}}(R)$ via Gaussian convolutions, are parametrized by a smearing convolution parameter $S$ and defined as [8, 9]
@@ II Smeared phase space distributions
with allowed parameter values $1\geq R$ and $R>S$; we see that the Sudarshan-Glauber distribution $P(x,p)={\tilde{\cal P}}(x,p;R=1)$ (whilst taking a suitable limit of $S\uparrow 1$) serves as a starting point of sorts.
Expression (1) connects all $S$-parametrized phase space distributions ${\tilde{\cal P}}(S)$. The ${\tilde{\cal P}}(S)$ are real-valued and with dropping values of $S$ increasingly more smeared convolutions of $P$. They include as specially named cases, the “exceedingly singular” [4] Sudarshan-Glauber $P$-distribution, as well as Wigner’s distribution $W\equiv{\tilde{\cal P}}(0)$ [13] and Husimi’s distribution $Q\equiv{\tilde{\cal P}}(-1)$ [14]. Notably, $W$ can have negative values in phase space but is a well-behaved function [50] and stands out as the only distribution ${\tilde{\cal P}}(S)$ whose projections exactly give the state $\varrho$}}’s quadrature probability densities [10]. $Q$ stands out as the distribution with the largest possible value of $S$ where the smearing is sufficient to guarantee that $Q(x,p)\geq 0$ throughout phase space for any quantum state $\varrho$}}.
But, whilst being well-behaved, neither $W$ nor $Q$ can serve as a function to certify that a state $\varrho$}} is nonclassical. $Q$ has lost the ability to develop negative values and also in the case of $W$ too much gaussian-convolution induced smearing has happened: $W(x,p)<0$ is sufficient to certify that a state $\varrho$}} is nonclassical but it is not necessary for a state to have regions with negative values $W$ to be nonclassical [1, 11].
@@
The explicit form of the family of $S$-parametrized certification functionals devised in Ref. [1] is
$\xi(T)={\tilde{\xi}}(S)={\tilde{\cal P}}(S)-4\pi(1-S)\;[{\tilde{\cal P}}(2S-1)]^{2},$ (2)
@@
For transparency, we now perform the switch-over to the positive parametrization T=1ST=1-S, $T>0$ in Eq. (1) whilst dropping the tilde symbol, both, in expressions for $\xi$ and $\cal P$}. Next, in the spirit of Eq. (2), we form the difference between a phase space distribution ${\cal P}(T)$ and its smoothened version, ${\cal P}(T+{\Delta T})$, where the latter is rescaled and remapped, to define the quantumness certification functional
${\cal S}(x,p;T,{\Delta T})=$ $\;{\cal P}\left(x,p;T\right)-N_{0}\left(T\right)\left(\frac{{\cal P}\left(x,p;T+{\Delta T}\right)}{N_{0}\left(T+{\Delta T}\right)}\right){}^{\frac{T+{\Delta T}}{T}}$ (3)
@@
The logic behind this construction is this: We want $\cal S$}, just like $\xi$, to map coherent states to the value zero. Since coherent states are displaced vacuum states $\rho=|0\rangle\langle 0|$ we use that ${\cal P}_{|0\rangle\langle 0|}(T)=N_{0}(T)\exp[-(x^{2}+p^{2})/T]$ with norm $N_{0}\left(T\right)=\tfrac{1}{\pi T}$. When mapped to ${\cal P}_{|0\rangle\langle 0|}(T+\Delta T)$, the peak height drops to $N_{0}(T+\Delta T)$ and the distribution becomes spread out. To compensate for the loss in height we divide by the norm, giving a peak with unit height. The exponentiation by the fraction $\tfrac{T+{\Delta T}}{T}$} undoes the spreading, a further rescaling by the factor $N_{0}(T)$ matches the twosummands with each other such that $\cal S$} identically vanishes for all pure coherent states: it is desirable for pure coherent states to have ${\cal S}(x,p;T,{\Delta T})\equiv 0$ since coherent states ‘live’ right on the dividing line between classical and quantum states.
We note that for $\cal S$}, just like for $\xi$, negative values somewhere in phase space for some values of $T$, certify nonclassical behaviour [1]. But especially for experimentally generated states, distributions less smoothened than Wigner’s distribution, ${\cal P}(x,p;1)$, cannot be reconstructed on account of having to use numerically unstable reconstruction schemes [10]. We therefore only consider Wigner distributions or those which are more smeared, ${\cal P}(x,p;T)$ with $T\geq 1$, as relevant to our discussion, also see Fig. 1.
We find for ${\Delta T}=T$ that ${\cal S}(T,T)=\xi(T)$, in other words, ${\cal S}(T,{\Delta T})$ generalizes $\xi(T)$, this is the main result of this work.
III $\xi(T)$ can be faithfully discriminating
@@
A particular strength of the certification function $\xi$ of Ref. [1] is that, without fail, it detects below-vacuum fluctuations in squeezed states. Irrespective of whether the states are pure (Fig. 1 of [1]) or impure squeezed (gaussian) states. This can be understood from the argument given above when deriving Eq. (3) for $\cal S$}. The exponentiation by the fraction $\tfrac{T+{\Delta T}}{T}$}, for Gaussians which decline faster than vacuum states, does not fully undo the previous spreading when mapping ${\cal P}\left(x,p;T\right)$ to ${\cal P}\left(x,p;T+{\Delta T}\right)$: upon forming their difference in $\cal S$} a region with negative values arises on the squeezed ‘narrow flanks’ of the phase space distribution (Fig. 1 of [1]), faithfully certifying nonclassical behaviour [11].
IV A case where first $\xi(T)$ and then ${\cal S}(T,{\Delta T})$ fails to be faithfully discriminating
In order to investigate a scenario where $\xi(T)$ fails to be faithfully discriminating, we consider a state with a mixed positive background distribution of a pair of two overlapping incoherently mixed (classical) vacuum states, $W_{0}$, mixed together with a single photon state, $W_{1}$, which carries a small weight $w_{1}\approx 1$%, namely:
$\!\!\rho(w_{1}$ $,s)=w_{1}W_{1}(x+\tfrac{1}{20},p)+(1-w_{1})$
$\times$ $\left(W_{0}(x+\tfrac{1}{20},p)\cos^{2}(s)+\sin^{2}(s)W_{0}(x-\tfrac{37}{20},p)\right)\!,\!$ (5)
with $s=\tfrac{\pi}{20}$. Fig. 1 shows that the certification functional $\xi$ of Ref. [1] is slightly less discriminating than the certification functional $\cal S$} of Eq. (3), but both fail when the single photon state’s weight $w_{1}$ drops further below 0.0122.
@@ IV A case where first $\xi(T)$ and then ${\cal S}(T,{\Delta T})$ fails to be faithfully discriminating
Figure 1: Here we study $\xi(T=1)$ and ${\cal S}(T=1,\Delta T)$ for Wigner’s distribution, ${\cal P}(x,p;T=1)$, since it is the borderline case experimentally accessible in state reconstruction (see remarks after Eq. (4)). We consider the weakly nonclassical state (5): for small values of $w_{1}\leq 0.0125$. The certification functional $\xi(x,p;1)$ of Ref. [1] (red band) fails to drop to negative values (below green zero-value sheets), i.e., it fails to detect the nonclassical contribution $w_{1}W_{1}$ in state (5) for the value $w_{1}=0.0122$ (bottom panels), whereas the certification functional ${\cal S}(x,p;T=1,\Delta T)$ of Eq. (3) still forms negative values. For slightly smaller values of $w_{1}<0.0122$ (not shown) $\cal S$} fails as well. In short, $\cal S$} can be more discriminating than $\xi$ but not faithfully so.

### v2

# arXiv:2512.23299v2[quant-ph] 02 Mar 2026

@@ Abstract
If the phase space-based Glauber-Sudarshan distribution, $P_{{{\varrho}}}$, has negative values the quantum state, $\varrho$}}, it describes is nonclassical. Due to $P$’s singular behaviour this simple criterion is impractical to use. Recent work [Bohmann and Agudelo, Phys. Rev. Lett. 124, 133601 (2020)] presented a general, sensitive, and noise-tolerant certification functional, $\xi[P]$, for the detection of non-classical behaviour of quantum states $P_{{{\varrho}}}$. There, it was shown that when this functional takes on negative values somewhere in phase space, $\xi[P](x,p)<0$, this is sufficient to certify the nonclassicality of a state. Here we give examples where this certification fails. We investigate states which are known to be nonclassical but the certification functions is non-negative, $\xi(x,p)\geq 0$, everywhere in phase space. We generalize $\xi$ giving it an appealing form which allows for improved certification. This way we generate a more sensitive family of certification functions. Yet, also these fail for very weakly nonclassical states, the question how to faithfully certify quantumness remains an open question.
Why Quantumness Certification Functionals?
It is desirable to be able to discriminate between quantum states on the one hand and classical states on the other. So far no simple approach that works for theoretical studies and experimentally reconstructed states alike has been found to perform this discrimination faithfully. All known approaches to discrimination are either unworkable in the general case, or they do not work unambiguously, or both [1, 2].
Here we generalize a recently devised, sensitive, quantumness certification functional, $\xi$ [3]. Our generalization provides insight into the structure of $\xi$ and makes it more sensitive still, yet, the goal of a universally faithful approach to quantum versus classical discrimination remains elusive.
In the 1960s it was established that the Glauber-Sudarshan phase space distribution, $P$ [4, 5, 6, 7, 8, 9, 10], can be used for the very definition of nonclassical behaviour [6, 7, 3]. If a state, $\varrho$}}, has a distribution $P_{{{\varrho}}}(x,p)$ with negative values somewhere at position $x$ and momentum $p$ in (cartesian) phase space  [^1] Instead of using complex coherent amplitudes to parameterize phase space we follow the convention for (quantum-mechanical) phase space of Ref. [10] using position $x$ and momentum $p$., then that state is nonclassical.
@@
Very useful alternatives to Glauber-Sudarshan’s $P$ distribution are Wigner’s, $W$ [15], and Husimi’s, $Q$ [16], distributions. Both are more well-behaved functions than $P$ is, but at the expense of not being able to certify that a state is nonclassical.
This situation has led to many different attempts to devise a reliable strategy for certification of the nonclassical nature of a state (see [17, 18, 19, 20, 21, 22, 23, 24, 3, 25, 26, 27] and the many references therein). Similarly, a plethora of measures of nonclassicality [28, 29, 30, 31, 32, 17, 33, 34, 35, 36, 37, 38, 39, 21, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 26, 52] have been devised. None of these approaches faithfully discriminate between classical and quantum states: faithful discrimination should always correctly detect and certify that a state is nonclassical, and it should certify only classical states as classical. This ability remains elusive.
Quantumness certification approaches devised so far, typically, mislabel some states that have weak quantum features as classical [43, 26]. In the context of phase space-based approaches, this is often due to the fact that mapping functions from the potentially very singular $P$-distributions to better behaved ones, such as Wigner and Husimi distributions, washes out the very features (where the distribution is negative-valued) we intend to detect [43], see Fig. 1.
Figure 1: $P$ (approximate), $W$ and $Q$ distributions for state (8).
Here we consider the recently devised phase space-based certification functional for single mode systems $\xi[{{\varrho}}](x,p)$, introduced by Bohmann and Agudelo in Ref. [3], in which they showed that all classical states obey $\xi[{{\varrho}}](x,p)\geq 0$. This is to be understood as the statement that the occurrence of negative values $\xi[{{\varrho}}](x,p)<0$ somewhere in phase space $(x,p)$ is sufficient to show that a state is nonclassical [3]. What is missing for $\xi$ of [3], however, is necessity: to be faithfully discriminating it would be necessary that whenever a state is nonclassical we find that $\xi(x,p)<0$ in phase space somewhere. Below, we will show that there are states $\rho$ which are (weakly) nonclassical and yet for these states $\xi_{\rho}\geq 0$, across phase space.
$\xi[{{\varrho}}](x,p)$ is a functional of fixed form independent of context such as the type of system or the type of state the system is in. $\xi[{{\varrho}}]$ applies to all states of one-dimensional continuous-variable systems, irrespective of whether the state is pure or mixed and whether it has small, large, or even macroscopic [53] extent, irrespective of context, hamiltonian, environment or measurement procedures [1], in this sense $\xi$ is universal, meaning it is applicable for all cases and to all states of single mode systems.
@@
Smeared phase space distributions
Smeared phase space distributions, ${\tilde{\cal P}}(S)$, expressed in terms of a reference distribution ${\tilde{\cal P}}(R)$ via Gaussian convolutions, are parameterized by a smearing convolution parameter $S$ and defined as [10, 11]
@@
with allowed parameter values $1\geq R$ and $R>S$; we see that the Glauber-Sudarshan distribution $P(x,p)={\tilde{\cal P}}(x,p;R=1)$ (whilst taking a suitable limit of $S\uparrow 1$) serves as a starting point of sorts.
Expression (1) connects all $S$-parameterized phase space distributions ${\tilde{\cal P}}(S)$. The ${\tilde{\cal P}}(S)$ are real-valued and with dropping values of $S$ increasingly more smeared convolutions of $P$. They include as specially named cases, the “exceedingly singular” [6] Glauber-Sudarshan $P$-distribution, as well as Wigner’s distribution $W\equiv{\tilde{\cal P}}(0)$ [15] and Husimi’s distribution $Q\equiv{\tilde{\cal P}}(-1)$ [16]. Notably, $W$ can have negative values in phase space but is a well-behaved function [54] and stands out as the only distribution ${\tilde{\cal P}}(S)$ whose projections exactly give the state $\varrho$}}’s quadrature probability densities [12]. $Q$ stands out as the distribution with the largest possible value of $S$, with just sufficient smearing to guarantee that $Q(x,p)\geq 0$ throughout phase space for any quantum state $\varrho$}}.
I The Certification Functionals
Neither $W$ nor $Q$ can serve as a function to certify that a state $\varrho$}} is nonclassical. $Q$ has lost the ability to develop negative values and also in the case of $W$ too much gaussian-convolution induced smearing has happened: $W(x,p)<0$ is sufficient to certify that a state $\varrho$}} is nonclassical but it is not necessary for a state to have regions with negative values $W$ to be nonclassical [3, 13].
@@ I The Certification Functionals
The explicit form of the family of $S$-parameterized certification functionals devised in Ref. [3] is
$\!\!\!\!\!\xi[{\cal P}](T)={\tilde{\xi}[\tilde{\cal P}]}(S)={\tilde{\cal P}}(S)-4\pi(1-S)\;[{\tilde{\cal P}}(2S-1)]^{2},\!\!\!\!$ (2)
@@
For transparency, we just performed the switch-over to the positive parameterization T=1ST=1-S, $T>0$ in Eq. (1) whilst dropping the tilde symbol, in both, in expressions for $\xi$, and for $\cal P$}. Next, in the spirit of Eq. (2), we form the difference between a phase space distribution ${\cal P}(T)$ and its smoothened version, ${\cal P}(T+{\Delta T})$, where the latter is rescaled and remapped, to define the quantumness certification functional
${\cal S}[{\cal P}](x,p;T,{\Delta T})=$ $\;{\cal P}\left(x,p;T\right)-N_{0}\left(T\right)\left(\frac{{\cal P}\left(x,p;T+{\Delta T}\right)}{N_{0}\left(T+{\Delta T}\right)}\right){}^{\frac{T+{\Delta T}}{T}}$ (3)
@@
The logic behind this construction is this: We want $\cal S$}, just like $\xi$, to map coherent states to the value zero. Since coherent states are displaced vacuum states $\rho=|0\rangle\langle 0|$ we use that ${\cal P}_{|0\rangle\langle 0|}(T)=N_{0}(T)\exp[-(x^{2}+p^{2})/T]$ with peak height $N_{0}\left(T\right)=\tfrac{1}{\pi T}$. When mapped to ${\cal P}_{|0\rangle\langle 0|}(T+\Delta T)$, this peak height drops to $N_{0}(T+\Delta T)$ and the distribution becomes spread out. To compensate for the loss in height we divide by the norm, giving us a peak with unit height. The exponentiation by the fraction $E={\tfrac{T+{\Delta T}}{T}}$ undoes the spreading, a further rescaling by the factor $N_{0}(T)$ matches the two summands with each other such that $\cal S$} identically vanishes for all pure coherent states [3]: it is desirable for pure coherent states to have ${\cal S}_{|0\rangle\langle 0|}(x,p;T,{\Delta T})\equiv 0$ since coherent states ‘live’ right on the dividing line between classical and quantum states.
${\cal S}(T,{\Delta T})$ generalizes $\xi(T)$
When ${\Delta T}=T$, ${\cal S}(T,T)=\xi(T)$, or, ${\cal S}(T,{\Delta T})$ generalizes $\xi(T)$, this is the main result of this work.
Our derivation also allows us to generalize to the multimode case, for $K$ cartesian phase space modes:
${\cal S}(T,{\Delta T})={\cal P}\left(T\right)-N^{K}_{0}\left(T\right)\left(\tfrac{{\cal P}\left(T+{\Delta T}\right)}{N^{K}_{0}\left(T+{\Delta T}\right)}\right){}^{\tfrac{T+{\Delta T}}{T}}$, (5)
This is the other main result of this work.
II $\xi(T)$ and $\cal S$} can always identify classical states
$\cal S$}, just like $\xi$, without fail, certifies classical states (incoherent mixtures of Glauber coherent states) as classical, since the nonlinear terms in Eq. (3) for a mixture between states ${\cal P}_{1}$ and ${\cal P}_{2}$, parameterized by ($0<w<1)$, form the positive concave difference function ${\cal S}[w{\cal P}_{1}+(1-w){\cal P}_{2}]-\left(w{\cal S}[{\cal P}_{1}]+(1-w){\cal S}[{\cal P}_{2}]\right)\geq 0$. This renders mixtures $\cal S$} more positive than the mixture’s individual contributions, as desired.
We now show this in two steps.
By construction ${\cal S}\equiv 0$ for pure coherent states. Since classical states are incoherent mixtures of Glauber coherent states, we first establish that a coherent state contributing to a mixture with below-unit weight $w$ ($0<w<1$) gives a positive contribution to ${\cal S}[w{\cal P}_{|0\rangle\langle 0|}]>0$: recall that Eq. (3) implies that ${\cal S}[w{\cal P}_{|0\rangle\langle 0|}](T,{\Delta T})$ $w\left({\cal P}_{|0\rangle\langle 0|}\left(T\right)-{w}^{\frac{{\Delta T}}{T}}N_{0}\left(T\right)\left(\frac{{\cal P}_{|0\rangle\langle 0|}\left(T+{\Delta T}\right)}{N_{0}\left(T+{\Delta T}\right)}\right){}^{\frac{T+{\Delta T}}{T}}\right)>w\,{\cal S}[{\cal P}_{|0\rangle\langle 0|}](T,{\Delta T})=0,$ since ${w}^{\frac{{\Delta T}}{T}}<1$.
III $\xi(T)$ and $\cal S$} are Concave functionals of the mixing of states
In this second step, we now generalize this result by showing, for all states, mixtures render $\cal S$} more positive than the mixture’s individual contributions. This reduces the telltale negative values of nonclassical states for mixtures.
$\cal S$} is strictly concave in $w$:
$\Delta{\cal S}(T,\Delta T)=\;$ ${\cal S}[w{\cal P}_{1}+(1-w){\cal P}_{2}]$
$-\Big(w{\cal S}[{\cal P}_{1}]+(1-w){\cal S}[{\cal P}_{2}]\Big)$ (6)
$\propto$ $-\left[w{\cal P}_{1}+(1-w){\cal P}_{2}\right]^{E}$
$+\left(w[{\cal P}_{1}]^{E}+(1-w)[{\cal P}_{2}]^{E}\right)>0;$ (7)
[…]
This concave mixedness relation is the second main result of this work.
IV $\xi(T)$ can be faithfully discriminating for some states
@@
A particular strength of the certification function $\xi$ of Ref. [3] is its ability, without fail, to detect squeezing of fluctuations below vacuum level in all gaussian states; irrespective of whether the states are pure (Fig. 1 of [3]) or impure squeezed states. This can be understood from the arguments given above when deriving Eq. (3) for $\cal S$}: For Gaussians which decline faster than vacuum states, the spreading when mapping ${\cal P}\left(T\right)$ to ${\cal P}\left(T+{\Delta T}\right)$ is not fully compensated for by the narrowing due to the exponentiation by $E={\tfrac{T+{\Delta T}}{T}}$. Upon forming the difference in $\cal S$} a region with negative values arises on the squeezed ‘narrow flanks’ of the phase space distribution (Fig. 1 of [3]), faithfully certifying nonclassical suppression of fluctuations [13] for all Gaussians.
V A case where first $\xi(T)$ and then ${\cal S}(T,{\Delta T})$ fails to be faithfully discriminating
@@
Figure 2: Here we study $\xi(T=1)$ and ${\cal S}(T=1,\Delta T)$ for Wigner’s distribution, ${\cal P}(T=1)$, since it is the borderline case experimentally accessible in state reconstruction (see the remarks at the beginning of this section). We consider the weakly nonclassical state (8): for small values of $w_{1}\leq 0.0125$. The certification functional $\xi(x,0;1)$ of Ref. [3] (red band) fails to drop to negative values (below green zero-value sheets), i.e., it fails to detect the nonclassical contribution $w_{1}W_{1}$ in state (8) for the value $w_{1}=0.0122$ (bottom panels), whereas the certification functional ${\cal S}(x,0;T=1,\Delta T)$ of Eq. (3) still forms negative values. For slightly smaller values of $w_{1}<0.0122$ (not shown) $\cal S$} fails as well. In short, $\cal S$} can be more discriminating than $\xi$ but not faithfully so.
We note that for $\cal S$}, just like for $\xi$, negative values somewhere in phase space for some values of $T$, certify nonclassical behaviour [3]. But especially for experimentally generated states, distributions less smoothened than Wigner’s distribution, ${\cal P}(x,p;1)$, cannot be reconstructed on account of having to use numerically unstable reconstruction schemes [12]. Hence, below, we mostly consider Wigner distributions or those which are more smeared, ${\cal P}(x,p;T)$ with $T\geq 1$, see Fig. 2.
In order to investigate a scenario where $\xi(T)$ fails to be faithfully discriminating, we consider a state with a mixed positive background distribution of a pair of two overlapping incoherently mixed (classical) vacuum states, $W_{0}$, mixed together with a single photon Fock state, $W_{1}$, which carries a small weight $w_{1}\approx 1$%, namely:
$\!\!\rho(w_{1}$ $,s)=w_{1}W_{1}(x,p)+(1-w_{1})$
$\times$ $\Big(W_{0}(x,p)\cos^{2}(s)+\sin^{2}(s)W_{0}(x-\tfrac{19}{40},p)\Big),\!$ (8)
with $s=\tfrac{\pi}{20}$, see Fig. 1. Fig. 2 shows that for small values of $\Delta T$ the certification functional ${\cal S}(T,\Delta T)$ (3) can be slightly more discriminating than ${\cal S}(T,T)=\xi(T)$ of Ref. [3], but both fail when the single photon state’s weight $w_{1}$ drops further, below 0.0122. This is due to inequality (7).